# Byzantine-resilient Federated Learning With Adaptivity to Data Heterogeneity

### v2

@@ I-A Related Works
[24] RANGE non-convex strongly-convex IID one step non-zero -
@@
As of the aggregation strategies in related literature, by leveraging median [16], trimmed mean [19], or iterative filtering [20], the associated FL system can tolerate a small fraction of Byzantine users. To be more robust to Byzantine attacks, some methods such as Krum [21], Federated Learning Over the Air (FLOA) [22], Robust Stochastic Aggregation (RSA) [23], and Robust Aggregating Normalized Gradient Method (RANGE) [24], have been proposed. Specifically, Krum selects a stochastic gradient as the global one, which has the shortest cumulative distance from a group of stochastic gradients that are most closely distributed, FLOA requires each user to normalize the associated local gradient to be a vector with zero mean and unit variance before offloading it to the central server, RSA penalizes the difference between local model parameters and global model parameters to isolate Byzantine users, and RANGE takes the coordinate-wise median of uploaded stochastic gradients as the aggregated one. Other strategies try to filter out the messages from Byzantine attackers before aggregation at the central server through various measures, e.g., [28] leverges semi verified mean estimation and [29] utilizes a multi-dimensional Huber loss function.
Although the above aggregation strategies can be resistant to Byzantine attacks at some extent, they may still fail to work in front of heavy Byzantine attacks: The attacks with the fraction of Byzantine users being large or the amplitude of uploaded vectors by Byzantine users being high. Recently, geometric median appears as a good candidate for aggregation because of its ability to naturally tolerate a larger fraction of strong Byzantine attacks [30]. Based on geometric median, some algorithms such as Robust Federated Aggregation (RFA) [25], Byzantine attack resilient distributed Stochastic Average Gradient Algorithm (Byrd-SAGA) [26], and Byzantine-RObust Aggregation with gradient Difference Compression And STochastic variance reduction (BROADCAST) [27], have been proposed. To be specific, geometric median is leveraged to aggregate the uploaded vectors for RFA, Byrd-SAGA, and BROADCAST. Differently, each user in RFA selects the trimmed mean of the model parameters over multiple local updating as the uploaded vector, while Byrd-SAGA and BROADCAST utilize the SAGA [31] method to generate the vectors for aggregation first and then upload them in a lossless or compressive way.

### v3

@@ I-A Related Works
[21] RANGE non-convex strongly-convex IID one step non-zero -
[22] - non-convex strongly-convex IID little data heterogeneity multiple step non-zero Sub-exponential distribution
[23] - strongly-convex IID multiple step non-zero Sub-exponential distribution
@@
As of the aggregation strategies in related literature, by leveraging median [16], trimmed mean [19], or iterative filtering [20], the associated FL system can tolerate a small fraction of Byzantine users. To be more robust to Byzantine attacks, some methods such as Robust Aggregating Normalized Gradient Method (RANGE) [21], a Huber Loss Minimization [22], semi verified mean estimation [23], Krum [24], Federated Learning Over the Air (FLOA) [25], and Robust Stochastic Aggregation (RSA) [26], have been proposed. Specifically, RANGE takes the coordinate-wise median of uploaded stochastic gradients as the aggregated one. [22] improve the trimmed mean on non-IID datasets by utilizing a multi-dimensional Huber loss function to aggregate uploaded messages. [23] leverges semi verified mean estimation to filter out the messages from Byzantine attackers before aggregation. Krum selects a stochastic gradient as the global one, which has the shortest cumulative distance from a group of stochastic gradients that are most closely distributed, FLOA requires each user to normalize the associated local gradient to be a vector with zero mean and unit variance before offloading it to the central server, RSA penalizes the difference between local model parameters and global model parameters to isolate Byzantine users.
Although the above aggregation strategies can be resistant to Byzantine attacks at some extent, they may still fail to work in front of heavy Byzantine attacks without identifying the Byzantine attacker: The attacks with the fraction of Byzantine users being large or the amplitude of uploaded vectors by Byzantine users being high. Recently, geometric median appears as a good candidate for aggregation because of its ability to naturally tolerate a larger fraction of strong Byzantine attacks [30]. Based on geometric median, some algorithms such as Robust Federated Aggregation (RFA) [27], Byzantine attack resilient distributed Stochastic Average Gradient Algorithm (Byrd-SAGA) [28], and Byzantine-RObust Aggregation with gradient Difference Compression And STochastic variance reduction (BROADCAST) [29], have been proposed. To be specific, geometric median is leveraged to aggregate the uploaded vectors for RFA, Byrd-SAGA, and BROADCAST. Differently, each user in RFA selects the trimmed mean of the model parameters over multiple local updating as the uploaded vector, while Byrd-SAGA and BROADCAST utilize the SAGA [31] method to generate the vectors for aggregation first and then upload them in a lossless or compressive way.